OpenStudy (anonymous):

Verify that the vector X is a solution of the given system: dx/dt = 3x - 4y dy / dt = 4x - 7y \[x = \left(\begin{matrix}1 \\ 2\end{matrix}\right) e ^{-5t}\]

OpenStudy (perl):

this is a verification/substitution problem , would you like help :)

OpenStudy (anonymous):

Yeah I would. I know the rule states that if the answer can be derived, than it is a solution vector, but I'm not too sure how to get to there.

OpenStudy (anonymous):

That I get, and we can multiply to the initial system, right?

OpenStudy (perl):

$$ { \bf \vec x } = \left(\begin{matrix}x \\ y\end{matrix}\right) = \left(\begin{matrix}1\cdot e ^{-5t} \\ 2 \cdot e ^{-5t}\end{matrix}\right) $$

OpenStudy (perl):

yes, then we set x(t) = e^{-5t} y(t) = 2e^{-5t} find dx/dt, dy/dt then plug in

OpenStudy (anonymous):

yeah so matrix multiplication, I would get 3e^-5t - 8 (e^-5t) for dx/dt and 4e^-5t -14e^-5t for dy/dt right?

OpenStudy (perl):

I got x(t) = e^{-5t} dx/dt = -5e^{-5t} y(t) = 2e^{-5t} dy/dt = 2(-5)e^{-5t}

OpenStudy (perl):

ok one moment, two steps behind you

OpenStudy (anonymous):

oh wait you derived it. my bad. I thought you did a matrix multiplication. Yeah I get the samething when I derive

OpenStudy (perl):

ok :)

OpenStudy (anonymous):

why would you derive right away and not make a matrix multiplication?

OpenStudy (perl):

this is a checking problem. you want to check that x(t) = e^{-5t}, y(t) = 2e^{-5t} satisfy the system of equations : dx/dt = 3x - 4y dy / dt = 4x - 7y

OpenStudy (perl):

the x(t), y(t) are the components of the vector solution x

OpenStudy (anonymous):

Oh I see, so because my system are derivatived, I need to do the same to my solutions?

OpenStudy (perl):

right

OpenStudy (perl):

for example i get dx/dt = -5 e^{-5t} 3x - 4y = 3e^{-5t} - 4 (2e^{-5t}) = (3-8) e^{-5t} =-5 e^{-5t} Therefore dx/dt = 3x-4y \( \Large \checkmark \)

OpenStudy (anonymous):

ohhh I see ,that makes sense, thank you ^^ So the derivative of the solution must be equal to the solution space?

OpenStudy (anonymous):

to the system* my bad

OpenStudy (perl):

right :)

OpenStudy (perl):

Are you studying linear algebra?

OpenStudy (anonymous):

No differential equations. I already studied linear algebra :)

OpenStudy (perl):

ok :)